Suppose the time spent by a randomly selected student at a campus computer lab has a gamma distribution with mean 20 minutes and variance 80 minutes.


1
Expert's answer
2022-03-01T16:00:40-0500

a. What are the values of "\u03b1" and "\u03b2"


"\mu =\alpha \beta=20"

"\sigma^2=\alpha \beta^2=80"

"\beta=\dfrac{80}{20}=4"

"\alpha=\dfrac{20}{4}=5"

b. What is the probability that a student uses the terminal for at most 24 min?


"P(X\le 24)=F(24;\alpha, \beta)=F(24/\beta;\alpha)"

"=F(24\/4;5)=F(6;5)=0.7149"